In the last quarter of​ 2007, a group of 64 mutual funds had a mean return of

2.4%

with a standard deviation of

6.1%

Consider the Normal model

​N(0.0240.024​,0.0610.061​)

for the returns of these mutual funds.

​a) What value represents the 40th percentile of these​ returns?

​b) What value represents the 99th​ percentile?

​c) What's the​ IQR, or interquartile​ range, of the quarterly returns for this group of​ funds?
Answer #1

(a)

Ans : 0.01 = 1%

(b)

Ans : 0.17 = 17%

(c)

IQR = Q3 - Q1 = 0.06+0.02 = 0.08 = 8%